Two balls are thrown simultaneously, A vertically upwards with a speed of 20m/s from the ground, and B vertically downwards from a height of 40m with same speed and along same line of motion. At what points do the balls meet?

Grade:11

1 Answers

Mayank Sharma
6 years ago
Let balls meet at a distance of xmeter from top of the tower at time t.for ball A u= -20m/s, S= -x meter, a=-9.8m/s^2,using eqn S = u*t+ 1/2a*t^2 -x= -20*t+1/2*(-9.8)*t^2 - eqn ifor ball B u= 20 m/s, S=(40-x)m, a=-9.8m/s^2using eqn S=u*t+1/2a*t^240-x=20*t+1/2*(-9.8)*t^2 - eqn iiputting -x from eqn i in eqn ii,40-20*t+1/2(-9.8)*t^2=20*t+1/2(-9.8)*t^240-20*t=20*t (linear eqn in t hence only one soln ie balls meet only once )40=40*t ,=> t=1sputting t=1s in eqn i we get-x=-20+1/2(-9.8)-x=-20-4.9-x=-24.9x=24.9mhence balls meet only once at t=1s, at a distance of 24.9m from the top of the tower. or at a height of 15.1m(40-24.9=15.1).Hope this helps :)